Solve by substitution
4x-12y=5
-x+3y=-1

Answer by fractalier(6550)

Well, let us solve the second one for x and substitute and see what we get...so from
4x - 12y = 5
-x + 3y = -1
we have
x = 3y + 1
and then
4(3y + 1) - 12y = 5
12y + 4 - 12y = 5
4 = 5
You then ask, "When does 4 = 5?"
Never of course.
Your conclusion is there is no solution. These lines are parallel.